OpenStudy (ccchristian1):

Variable x is 7 more than variable y. Variable x is also 1 less than y. Which of the following pairs of equations best models the relationship between x and y?

OpenStudy (ccchristian1):

1. x = 1y x = y + 7 2. x = y + 7 x = y − 1 3. y = x + 8 y = x − 1 4. y = 7x y = x + 1

OpenStudy (ccchristian1):

@triciaal

OpenStudy (ccchristian1):

@TheWeeknd16

OpenStudy (theweeknd16):

B.

OpenStudy (theweeknd16):

@ccchristian1

OpenStudy (triciaal):